OpenStudy (anonymous):

The graph of y = ax^2 + bx + c is a parabola that opens up and has a vertex at (-2, 5). What is the solution set of the related equation 0 = ax^2 + bx + c? a) Ø b) {-2} c) {5}

OpenStudy (cwrw238):

(x + 2)^2 + 5 = 0 is the equation in vertex form can you continue from here?

OpenStudy (anonymous):

-2?

OpenStudy (cwrw238):

expanding x^2 + 4x + 4 + 5 = 0 x^2 + 4x + 9 = 0 there are no real solutions to the so a) is correct

OpenStudy (anonymous):

oh... thanks!